### v1

# A note on higher coherence of graphs of groups

Abstract.
For $n\in\mathbb{N}$, a group is called $n$-coherent if every subgroup of type $\mathsf{F}_{n}$ is of type $\mathsf{F}_{n+1}$. For $n\geq 1$, we observe that graphs of groups with $n$-coherent vertex groups and virtually poly-cyclic edge groups are $n$-coherent. We deduce the $n$-coherence of certain right-angled Artin groups.
Key words and phrases:
Higher coherence, graphs of groups, right-angled Artin groups
2020 Mathematics Subject Classification:
57M07, 20F36
1. Introduction
A group is said to be coherent if every finitely generated subgroup is finitely presented. This notion has many connections, e.g., to low-dimensional topology, geometric group theory, graph theory, and ring theory [Wis20]. The subject of this note is a higher version of coherence.
For $n\in\mathbb{N}$, a group $G$ is of type $\mathsf{F}_{n}$ (resp. of type $\mathsf{F}_{\infty}$) if there exists a CW-model for the classifying space BGBG with finite $n$-skeleton (resp. of finite type). Every group is of type $\mathsf{F}_{0}$. For a group being of type $\mathsf{F}_{1}$ (resp. of type $\mathsf{F}_{2}$) is equivalent to being finitely generated (resp. finitely presented).
Definition $1.$
Let $n,m\in\mathbb{N}\cup\{\infty\}$ with $n<m$. We say that a group $G$ is $(n,m)$-coherent if every subgroup of $G$ that is of type $\mathsf{F}_{n}$ is of type $\mathsf{F}_{m}$.
Definition 1.1 generalises the usual notions: $(0,1)$-coherence is being Noetherian (also called slender), $(1,2)$-coherence is coherence, and $(n,n+1)$-coherence is recently called $n$-coherence [KV23, Vid, KV25]. Clearly, $(n,m)$-coherence implies $(n^{\prime},m^{\prime})$-coherence for $n\leq n^{\prime}<m^{\prime}\leq m$ and is preserved under taking subgroups. Moreover, $(n,m)$-coherence passes to finite index overgroups and hence is invariant under commensurability. Virtually poly-cyclic groups are $(0,\infty)$-coherent. Groups of geometric dimension $\leq n$ are $(n,\infty)$-coherent. The direct product of free groups $(F_{2})^{k}$ is $(n,n+1)$-coherent if and only if $n\geq k$.
The following is a classical combination theorem for coherence of graphs of groups.
Theorem 1.2 (Karrass–Solitar [KS70]).
Let $G$ be the fundamental group of a graph of groups whose vertex groups are $(1,2)$-coherent and whose edge groups are $(0,1)$-coherent. Then $G$ is $(1,2)$-coherent.
We prove a combination theorem for higher coherence of graphs of groups. We use the convention that $\infty-1=\infty$.
Theorem 1.3.
Let $n,m\in\mathbb{N}\cup\{\infty\}$ with $1\leq n<m$. Let $G$ be the fundamental group of a graph of groups whose vertex groups are $(n,m)$-coherent and whose edge groups are
$\begin{cases}(0,m-1)\text{-coherent}&\text{if }n=1;\\ (1,m-1)\text{-coherent}&\text{if }n\geq 2.\end{cases}$
Then $G$ is $(n,m)$-coherent.
Theorem 1.2 was crucial to characterise the coherent right-angled Artin groups. A simplicial complex is flag if every complete subgraph in its 1-skeleton spans a simplex. Associated to a finite flag simplicial complex $L$ is the right-angled Artin group $A_{L}$ that has as generating set the vertices of $L$, subject only to the relations that two generators commute if the corresponding vertices are adjacent in $L$ [Kob22]. A graph is chordal if it does not contain a cycle of length $\geq 4$ as a full subgraph. It is a classical result of Dirac [Dir61] that in chordal graphs, minimal separating sets are cliques. It follows that a flag simplicial complex with chordal 11-skeleton is a “tree of simplices” in the following sense.
Definition 1.4.
We define $\mathcal{T}_{1}$ to be the smallest class of finite flag simplicial complexes that satisfies the following:
  1. ()
    All simplices (of all dimensions) lie in $\mathcal{T}_{1}$;
  2. ()
    If $L_{1},L_{2}\in\mathcal{T}_{1}$ and $L_{0}$ is a full subcomplex of both $L_{1}$ and $L_{2}$ that is a simplex or is empty, then the gluing $L_{1}\cup_{L_{0}}L_{2}$ lies in $\mathcal{T}_{1}$.
Theorem 1.5 (Droms [Dro87]).
Let $L$ be a finite flag simplicial complex. Then the following are equivalent:
  1. ()
    The group $A_{L}$ is $(1,\infty)$-coherent;
  2. ()
    The group $A_{L}$ is $(1,2)$-coherent;
  3. ()
    The 11-skeleton of $L$ is a chordal graph;
  4. ()
    The simplicial complex $L$ lies in the class $\mathcal{T}_{1}$.
We consider similar classes of simplicial complexes that allow for more vertex spaces and for edge spaces in $\mathcal{T}_{1}$.
Definition 1.6.
Let $n\in\mathbb{N}$ with $n\geq 2$. We define $\mathcal{T}_{n}$ to be the smallest class of finite flag simplicial complexes that satisfies the following:
  1. ()
    All simplices (of all dimensions) lie in $\mathcal{T}_{n}$;
  2. ()
    All finite flag simplicial complexes of dimension $\leq n-1$ lie in $\mathcal{T}_{n}$;
  3. ()
    If $L_{1},L_{2}\in\mathcal{T}_{n}$ and $L_{0}$ is a full subcomplex of both $L_{1}$ and $L_{2}$ that lies in $\mathcal{T}_{1}$ or is empty, then the gluing $L_{1}\cup_{L_{0}}L_{2}$ lies in $\mathcal{T}_{n}$;
  4. ()
    If $L\in\mathcal{T}_{n}$, then the cone $L\ast\Delta^{0}$ lies in $\mathcal{T}_{n}$.
As an application of Theorem 1.3, we deduce the higher coherence of certain right-angled Artin groups.
Theorem 1.7.
Let $n\in\mathbb{N}$ with $n\geq 2$ and let $L$ be a finite flag simplicial complex. If $L\in\mathcal{T}_{n}$, then the right-angled Artin group $A_{L}$ is $(n,\infty)$-coherent.
We do not expect the converse of Theorem 1.7 to hold. This leaves open a characterisation of higher coherence for right-angled Artin groups.
2. Inheritance of higher coherence
We recall some well-known inheritance results for finiteness properties [Bie81, Bro82, Geo08, HW21].
Theorem 2.1.
Let $G$ be the fundamental group of a finite graph of groups and let $n\in\mathbb{N}\cup\{\infty\}$. The following hold:
  1. ()
    If all vertex groups are of type $\mathsf{F}_{n}$ and all edge groups are of type $\mathsf{F}_{n-1}$, then $G$ is of type $\mathsf{F}_{n}$;
  2. ()
    If $G$ is of type $\mathsf{F}_{n}$ and all edge groups are of type $\mathsf{F}_{n}$, then all vertex groups are of type $\mathsf{F}_{n}$.
By Bass–Serre theory, there is a correspondence between graphs of groups and groups acting on trees (without inversions), under which vertex and edge groups correspond to vertex and edge stabilisers, respectively, of orbit representatives.
Theorem 2.2 ([GL17, Corollary 2.19]).
Let $G$ be a group and let $T$ be a cocompact $G$-tree. If $G$ is finitely presented, then there exists a cocompact $G$-tree $S$ with finitely generated edge stabilisers together with a $G$-map $S\to T$ that sends vertices to vertices and edges to non-trivial edge paths. In particular, the vertex and edge stabilisers of $S$ are subgroups of vertex and edge stabilisers of $T$, respectively.
Theorem 2.3.
Let $1\to N\to G\to Q\to 1$ be a group extension and let $n\in\mathbb{N}\cup\{\infty\}$. The following hold:
  1. ()
    If $Q$ is of type $\mathsf{F}_{n}$ and NN is of type $\mathsf{F}_{n}$, then $G$ is of type $\mathsf{F}_{n}$;
  2. ()
    If $G$ is of type $\mathsf{F}_{n}$ and NN is of type $\mathsf{F}_{n-1}$, then $Q$ is of type $\mathsf{F}_{n}$.
These inheritance results for finiteness properties give rise to corresponding inheritance results for higher coherence.
Proof of Theorem 1.3.
Let $H$ be a subgroup of $G$ that is of type $\mathsf{F}_{n}$. We show that $H$ is of type $\mathsf{F}_{m}$. Consider the Bass–Serre tree $T$ associated to $G$ as an $H$-tree by restriction. Since the stabilisers of the $H$-tree $T$ are subgroups of the stabilisers of the $G$-tree $T$, they inherit the corresponding coherence properties. Since $H$ is finitely generated, there exists a cocompact $H$-subtree $T^{\prime}$ of $T$. In particular, the vertex and edge stabilisers of the $H$-tree $T^{\prime}$ are subgroups of vertex and edge stabilisers of the $H$-tree $T$, respectively.
We treat the two cases $n=1$ and $n\geq 2$ separately. First, let $n=1$. Since the edge stabilisers of $T$ are (0,m1)(0,m-1)-coherent, the edge stabilisers of $T^{\prime}$ are of type $\mathsf{F}_{m-1}$ and, in particular, of type $\mathsf{F}_{1}$. By Theorem 2.1 ()2.1, the vertex stabilisers of $T^{\prime}$ are of type $\mathsf{F}_{1}$. Since the vertex stabilisers of $T$ are $(1,m)$-coherent, the vertex stabilisers of $T^{\prime}$ are of type $\mathsf{F}_{m}$. By Theorem 2.1 ()2.1, it follows that $H$ is of type $\mathsf{F}_{m}$.
Second, let $n\geq 2$. By Theorem 2.2, there exists a cocompact $H$-tree $S$ whose edge stabilisers are of type $\mathsf{F}_{1}$ and whose vertex and edge stabilisers are subgroups of vertex and edge stabilisers of $T^{\prime}$, respectively. Since the edge stabilisers of $T$ are (1,m1)(1,m-1)-coherent, the finitely generated edge stabilisers of $S$ are of type $\mathsf{F}_{m-1}$ and, in particular, of type $\mathsf{F}_{n}$. By Theorem 2.1 ()2.1, the vertex stabilisers of $S$ are of type $\mathsf{F}_{n}$. Since the vertex stabilisers of $T$ are $(n,m)$-coherent, the vertex stabilisers of $S$ are of type $\mathsf{F}_{m}$. By Theorem 2.1 ()2.1, it follows that $H$ is of type $\mathsf{F}_{m}$. ∎
Lemma 2.4.
Let $1\to N\to G\to Q\to 1$ be a group extension and let $n,m\in\mathbb{N}\cup\{\infty\}$ with $n<m$. The following hold:
  1. ()
    If $Q$ is $(n,m)$-coherent and NN is $(0,m)$-coherent, then $G$ is $(n,m)$-coherent;
  2. ()
    If $G$ is $(n,m)$-coherent and NN is of type $\mathsf{F}_{m-1}$, then $Q$ is $(n,m)$-coherent.
Proof.
Let $p\colon G\to Q$ denote the quotient map. For part (i), let $H$ be a subgroup of $G$ that is of type $\mathsf{F}_{n}$. Consider the group extension
$1\to N\cap H\to H\to p(H)\to 1.$
Since NN is $(0,m)$-coherent, the group $N\cap H$ is of type $\mathsf{F}_{m}$ and, in particular, of type $\mathsf{F}_{n-1}$. By Theorem 2.3 ()2.3, the group $p(H)$ is of type $\mathsf{F}_{n}$. Since $Q$ is $(n,m)$-coherent, the group $p(H)$ is of type $\mathsf{F}_{m}$. By Theorem 2.3 ()2.3, it follows that $H$ is of type $\mathsf{F}_{m}$.
For part (ii), let $K$ be a subgroup of $Q$ that is of type $\mathsf{F}_{n}$. Similarly, consider the group extension $1\to N\to p^{-1}(K)\to K\to 1$ and apply Theorem 2.3 ()2.3 and ()2.3. ∎
Remark 2.5.
In Theorem 1.3 for $n=1$, instead of assuming that the edge groups are (0,m1)(0,m-1)-coherent, it suffices that the intersection of any subgroup of $G$ that is of type $\mathsf{F}_{n}$ with any edge group is of type $\mathsf{F}_{m-1}$. Similarly, in Lemma 2.4 ()2.4, instead of assuming that NN is $(0,m)$-coherent, it suffices that the intersection of any subgroup of $G$ that is of type $\mathsf{F}_{n}$ with NN is of type $\mathsf{F}_{m}$.
3. Higher coherence of some right-angled Artin groups
We apply the inheritance results above to deduce the higher coherence of certain right-angled Artin groups.
Proof of Theorem 1.7.
We show that $(n,\infty)$-coherence is satisfied by the right-angled Artin groups associated to the flag simplicial complexes in Definition 1.6 ()1.6 and ()1.6 and is preserved under the operations in Definition 1.6 ()1.6 and ()1.6.
(i) If $L$ is a simplex, then the group $A_{L}$ is finitely generated free abelian and hence $(0,\infty)$-coherent and, in particular, $(n,\infty)$-coherent.
(ii) If $L$ is of dimension $\leq n-1$, then the group $A_{L}$ is of geometric dimension $\leq n$ and hence $(n,\infty)$-coherent [Geo08, Proposition 7.2.13].
(iii) For a gluing $L\coloneqq L_{1}\cup_{L_{0}}L_{2}$ with $L_{1},L_{2}\in\mathcal{T}_{n}$ and $L_{0}\in\mathcal{T}_{1}$ or $L=\emptyset$, the group $A_{L}$ splits as an amalgamated product $A_{L_{1}}\ast_{A_{L_{0}}}A_{L_{2}}$. (Here $A_{\emptyset}$ is the trivial group.) The group $A_{L_{0}}$ is $(1,\infty)$-coherent by Theorem 1.5. By induction on the number of vertices, we may assume that the groups $A_{L_{1}}$ and $A_{L_{2}}$ are $(n,\infty)$-coherent. Then the group $A_{L}$ is $(n,\infty)$-coherent by Theorem 1.3.
(iv) For a cone $CL\coloneqq L\ast\Delta^{0}$ with $L\in\mathcal{T}_{n}$, the group $A_{CL}$ splits as a direct product $A_{L}\times\mathbb{Z}$. By induction on the number of vertices, we may assume that the group $A_{L}$ is $(n,\infty)$-coherent. Then the group $A_{CL}$ is $(n,\infty)$-coherent by Lemma 2.4 ()2.4. ∎
We note that if a right-angled Artin group $A_{L}$ splits as an amalgamated product over a non-trivial $(1,2)$-coherent group, then there exists a separating full subcomplex of $L$ that lies in $\mathcal{T}_{1}$ [Hul21].
Remark 3.1.
Right-angled Artin groups are so-called graph products of $\mathbb{Z}$. Theorem 1.7 holds more generally for graph products of $(0,\infty)$-coherent groups when Definition 1.6 ()1.6 is modified with respect to their geometric dimension. The $(1,2)$-coherence of graph products was studied in [Var19].
Right-angled Artin groups can have subgroups with exotic finiteness properties, which are obvious obstructions to higher coherence. Let $L$ be a finite flag simplicial complex. The Bestvina–Brady group $H_{L}$ is defined as the kernel of the group homomorphism $A_{L}\to\mathbb{Z}$ that sends every standard generator of $A_{L}$ to $1\in\mathbb{Z}$. The finiteness properties of the group $H_{L}$ are determined by the higher connectivity of $L$ as follows.
Theorem 3.2 (Bestvina–Brady [BB97]).
Let $L$ be a finite flag simplicial complex and let $n\in\mathbb{N}$ with $n\geq 1$. Then the group $H_{L}$ is of type $\mathsf{F}_{n}$ if and only if $L$ is (n1)(n-1)-connected.
Corollary $3.$
Let $L$ be a finite flag simplicial complex and let $n\in\mathbb{N}$ with $n\geq 1$. If the right-angled Artin group $A_{L}$ is $(n,n+1)$-coherent, then every full subcomplex of $L$ that is (n1)(n-1)-connected is $n$-connected.
For $n=1$, the converse of Corollary 3.3 holds by Theorem 1.5. For $n\geq 2$, we do not know if the converse of Corollary 3.3 holds.
Example 3.4.
Let $L$ be the flag simplicial complex spanned by the graph in Figure 1. The right-angled Artin group $A_{L}$ is isomorphic to $F_{2}\times F_{2}\times(\mathbb{Z}^{2}\ast\mathbb{Z})$. The group $A_{L}$ is not $(1,2)$-coherent by Theorem 1.5 because $L$ contains a cycle of length 44 as a full subcomplex. The group $A_{L}$ is also not $(2,3)$-coherent by Corollary 3.3 because $L$ is 11-connected and not 22-connected. On the other hand, the group $A_{L}$ is $(4,\infty)$-coherent by Theorem 1.7 because $L$ is 33-dimensional and hence lies in the class $\mathcal{T}_{4}$. We do not know if the group $A_{L}$ is $(3,\infty)$-coherent, although we suspect it to be. Note that $L$ does not lie in the class $\mathcal{T}_{3}$, while every full subcomplex of $L$ that is 22-connected is 33-connected.
$\bullet\bullet\bullet\bullet\bullet\bullet\bullet$
Figure 1.
Remark 3.5.
There is an analogous notion of higher homological coherence [KV23, Fis]. Let $R$ be a ring and let $n\in\mathbb{N}$. A group $G$ is of type $\mathsf{FP}_{n}(R)$ (resp. of type $\mathsf{FP}_{\infty}(R)$) if the trivial RGRG-module $R$ admits a projective resolution that is finitely generated in degrees $\leq n$ (resp. in all degrees).
Let $n,m\in\mathbb{N}\cup\{\infty\}$ with $n<m$. We say that a group $G$ is homologically $(n,m)$-coherent over $R$ if every subgroup of $G$ that is of type $\mathsf{FP}_{n}(R)$ is of type $\mathsf{FP}_{m}(R)$. For $n\geq 2$, homological $(n,m)$-coherence over $\mathbb{Z}$ implies $(n,m)$-coherence because for $n\geq 2$, a group is of type $\mathsf{F}_{n}$ if and only if it is of type $\mathsf{F}_{2}$ and of type $\mathsf{FP}_{n}(\mathbb{Z})$. Inheritance results similar to the ones proved above hold also for higher homological coherence.
Groups of cohomological dimension $n$ over $R$ are homologically $(n,\infty)$-coherent over $R$ [Bro82, Proposition VIII.6.1]. Over the field $\mathbb{Q}$ (and suitably also over other fields), this can be improved if the top-dimensional $\ell^{2}$-Betti number vanishes [Fis, Proposition 2.3]: If $G$ is a locally indicable group with $\operatorname{cd}_{\mathbb{Q}}(G)=n$ and $b^{(2)}_{n}(G)=0$, then $G$ is homologically $(n-1,\infty)$-coherent over $\mathbb{Q}$.
Let $L$ be the flag simplicial complex spanned by the graph in Figure 1. The right-angled Artin group $A_{L}$ satisfies $\operatorname{cd}_{\mathbb{Q}}(A_{L})=4$ and by [DL03]
$b^{(2)}_{4}(A_{L})=\dim_{\mathbb{Q}}H_{3}(L;\mathbb{Q})=0.$
Hence $A_{L}$ is homologically $(3,\infty)$-coherent over $\mathbb{Q}$ (and, in fact, over every field [FHL24, Proposition 3.15][Fis, Proposition 2.3]).
Acknowledgements.
We thank Sam Fisher and Marco Linton for helpful comments. The first author was partially supported by the SFB 1085 Higher Invariants (Universität Regensburg, funded by the DFG).

### v2

@@
 2020 Mathematics Subject Classification:
-57M07, 20F36
+57M07, 20J05, 20F36
 1. Introduction
@@
 For $n\in\mathbb{N}$, a group $G$ is of type $\mathsf{F}_{n}$ (resp.
-of type $\mathsf{F}_{\infty}$) if there exists a CW-model for the classifying space BGBG with finite $n$-skeleton (resp.
+of type $\mathsf{F}_{\infty}$) if there exists a CW-model for the classifying space $BG$ with finite $n$-skeleton (resp.
 of finite type).
@@
 finitely presented).
-Definition $1.$
+Definition 1.1.
 Let $n,m\in\mathbb{N}\cup\{\infty\}$ with $n<m$. We say that a group $G$ is $(n,m)$-coherent if every subgroup of $G$ that is of type $\mathsf{F}_{n}$ is of type $\mathsf{F}_{m}$.
@@
 It is a classical result of Dirac [Dir61] that in chordal graphs, minimal separating sets are cliques.
-It follows that a flag simplicial complex with chordal 11-skeleton is a “tree of simplices” in the following sense.
+It follows that a flag simplicial complex with chordal $1$-skeleton is a “tree of simplices” in the following sense.
 Definition 1.4.
@@
   3. ()
-    The 11-skeleton of $L$ is a chordal graph;
+    The $1$-skeleton of $L$ is a chordal graph;
   4. ()
@@
     All finite flag simplicial complexes of dimension $\leq n-1$ lie in $\mathcal{T}_{n}$;
+  3. ()
+    All finite flag simplicial complexes $L$ of dimension $n$ with $H_{n}(L;\mathbb{Q})=0$ lie in $\mathcal{T}_{n}$;
   4. ()
@@
   1. ()
-    If $Q$ is of type $\mathsf{F}_{n}$ and NN is of type $\mathsf{F}_{n}$, then $G$ is of type $\mathsf{F}_{n}$;
+    If $Q$ is of type $\mathsf{F}_{n}$ and $N$ is of type $\mathsf{F}_{n}$, then $G$ is of type $\mathsf{F}_{n}$;
   2. ()
-    If $G$ is of type $\mathsf{F}_{n}$ and NN is of type $\mathsf{F}_{n-1}$, then $Q$ is of type $\mathsf{F}_{n}$.
+    If $G$ is of type $\mathsf{F}_{n}$ and $N$ is of type $\mathsf{F}_{n-1}$, then $Q$ is of type $\mathsf{F}_{n}$.
 These inheritance results for finiteness properties give rise to corresponding inheritance results for higher coherence.
@@
   1. ()
-    If $Q$ is $(n,m)$-coherent and NN is $(0,m)$-coherent, then $G$ is $(n,m)$-coherent;
+    If $Q$ is $(n,m)$-coherent and $N$ is $(0,m)$-coherent, then $G$ is $(n,m)$-coherent;
   2. ()
-    If $G$ is $(n,m)$-coherent and NN is of type $\mathsf{F}_{m-1}$, then $Q$ is $(n,m)$-coherent.
+    If $G$ is $(n,m)$-coherent and $N$ is of type $\mathsf{F}_{m-1}$, then $Q$ is $(n,m)$-coherent.
 Proof.
@@
 $1\to N\cap H\to H\to p(H)\to 1.$
-Since NN is $(0,m)$-coherent, the group $N\cap H$ is of type $\mathsf{F}_{m}$ and, in particular, of type $\mathsf{F}_{n-1}$.
+Since $N$ is $(0,m)$-coherent, the group $N\cap H$ is of type $\mathsf{F}_{m}$ and, in particular, of type $\mathsf{F}_{n-1}$.
 By Theorem 2.3 ()2.3, the group $p(H)$ is of type $\mathsf{F}_{n}$.
@@
 In Theorem 1.3 for $n=1$, instead of assuming that the edge groups are (0,m1)(0,m-1)-coherent, it suffices that the intersection of any subgroup of $G$ that is of type $\mathsf{F}_{n}$ with any edge group is of type $\mathsf{F}_{m-1}$.
-Similarly, in Lemma 2.4 ()2.4, instead of assuming that NN is $(0,m)$-coherent, it suffices that the intersection of any subgroup of $G$ that is of type $\mathsf{F}_{n}$ with NN is of type $\mathsf{F}_{m}$.
+Similarly, in Lemma 2.4 ()2.4, instead of assuming that $N$ is $(0,m)$-coherent, it suffices that the intersection of any subgroup of $G$ that is of type $\mathsf{F}_{n}$ with $N$ is of type $\mathsf{F}_{m}$.
+3. Higher homological coherence
+There is an analogous notion of higher homological coherence [KV23, Fis].
+Let $R$ be a ring and let $n\in\mathbb{N}$.
+A group $G$ is of type $\mathsf{FP}_{n}(R)$ (resp.
+of type $\mathsf{FP}_{\infty}(R)$) if the trivial $RG$-module $R$ admits a projective resolution that is finitely generated in degrees $\leq n$ (resp.
+in all degrees).
+Definition 3.1.
+Let $R$ be a ring and let $n,m\in\mathbb{N}\cup\{\infty\}$ with $n<m$.
+We say that a group $G$ is homologically $(n,m)$-coherent over $R$ if every subgroup of $G$ that is of type $\mathsf{FP}_{n}(R)$ is of type $\mathsf{FP}_{m}(R)$.
+For $n\geq 2$, homological $(n,m)$-coherence over $\mathbb{Z}$ implies $(n,m)$-coherence because for $n\geq 2$, a group is of type $\mathsf{F}_{n}$ if and only if it is of type $\mathsf{F}_{2}$ and of type $\mathsf{FP}_{n}(\mathbb{Z})$.
+Inheritance results similar to Theorem 1.3 and Lemma 2.4 hold also for higher homological coherence.
+Groups of cohomological dimension $n$ over $R$ are homologically $(n,\infty)$-coherent over $R$ [Bro82, Proposition VIII.6.1].
+Over the ring $\mathbb{Z}$ (and suitably also over fields), this can be improved if the top-dimensional $\ell^{2}$-Betti number vanishes.
+Theorem 3.2 (Fisher, Jaikin-Zapirain–Linton).
+Let $n\in\mathbb{N}$ with $n\geq 1$ and let $G$ be a locally indicable group with $\operatorname{cd}_{\mathbb{Z}}(G)=n$.
+If $b^{(2)}_{n}(G)=0$, then $G$ is homologically $(n-1,\infty)$-coherent over $\mathbb{Z}$.
+Proof.
+The proof is analogous to that of [Fis, Proposition 2.3], see also [JZL25, Theorem 3.10]. ∎
 4. Higher coherence of some right-angled Artin groups
@@
 Proof of Theorem 1.7.
-We show that $(n,\infty)$-coherence is satisfied by the right-angled Artin groups associated to the flag simplicial complexes in Definition 1.6 ()1.6 and ()1.6 and is preserved under the operations in Definition 1.6 ()1.6 and ()1.6.
+We show that $(n,\infty)$-coherence is satisfied by the right-angled Artin groups associated to the flag simplicial complexes in Definition 1.6 ()1.6, ()1.6, and ()1.6 and is preserved under the operations in Definition 1.6 ()1.6 and ()1.6.
 (i) If $L$ is a simplex, then the group $A_{L}$ is finitely generated free abelian and hence $(0,\infty)$-coherent and, in particular, $(n,\infty)$-coherent.
 (ii) If $L$ is of dimension $\leq n-1$, then the group $A_{L}$ is of geometric dimension $\leq n$ and hence $(n,\infty)$-coherent [Geo08, Proposition 7.2.13].
-(iii) For a gluing $L\coloneqq L_{1}\cup_{L_{0}}L_{2}$ with $L_{1},L_{2}\in\mathcal{T}_{n}$ and $L_{0}\in\mathcal{T}_{1}$ or $L=\emptyset$, the group $A_{L}$ splits as an amalgamated product $A_{L_{1}}\ast_{A_{L_{0}}}A_{L_{2}}$.
+(iii) If $L$ is of dimension $n$ with $H_{n}(L;\mathbb{Q})=0$, then the group $A_{L}$ satisfies $\operatorname{cd}_{\mathbb{Z}}(A_{L})=n+1$ and $b^{(2)}_{n+1}(A_{L})=\dim_{\mathbb{Q}}H_{n}(L;\mathbb{Q})=0$ by [DL03].
+By Theorem 3.2, the group $A_{L}$ is homologically $(n,\infty)$-coherent over $\mathbb{Z}$ and, since $n\geq 2$, hence $(n,\infty)$-coherent.
+(iv) For a gluing $L\coloneqq L_{1}\cup_{L_{0}}L_{2}$ with $L_{1},L_{2}\in\mathcal{T}_{n}$ and $L_{0}\in\mathcal{T}_{1}$ or $L_{0}=\emptyset$, the group $A_{L}$ splits as an amalgamated product $A_{L_{1}}\ast_{A_{L_{0}}}A_{L_{2}}$.
 (Here $A_{\emptyset}$ is the trivial group.) The group $A_{L_{0}}$ is $(1,\infty)$-coherent by Theorem 1.5.
@@
 Then the group $A_{L}$ is $(n,\infty)$-coherent by Theorem 1.3.
-(iv) For a cone $CL\coloneqq L\ast\Delta^{0}$ with $L\in\mathcal{T}_{n}$, the group $A_{CL}$ splits as a direct product $A_{L}\times\mathbb{Z}$.
+(v) For a cone $CL\coloneqq L\ast\Delta^{0}$ with $L\in\mathcal{T}_{n}$, the group $A_{CL}$ splits as a direct product $A_{L}\times\mathbb{Z}$.
 By induction on the number of vertices, we may assume that the group $A_{L}$ is $(n,\infty)$-coherent.
@@
 Right-angled Artin groups are so-called graph products of $\mathbb{Z}$.
-Theorem 1.7 holds more generally for graph products of $(0,\infty)$-coherent groups when Definition 1.6 ()1.6 is modified with respect to their geometric dimension.
+Theorem 1.7 holds more generally for graph products of $(0,\infty)$-coherent groups when Definition 1.6 ()1.6 is modified suitably.
 The $(1,2)$-coherence of graph products was studied in [Var19].
+The $\ell^{2}$-Betti numbers and agrarian Betti numbers of graph products were computed in [DO12, FHL24].
 Right-angled Artin groups can have subgroups with exotic finiteness properties, which are obvious obstructions to higher coherence.
@@
 Let $L$ be a finite flag simplicial complex and let $n\in\mathbb{N}$ with $n\geq 1$. Then the group $H_{L}$ is of type $\mathsf{F}_{n}$ if and only if $L$ is (n1)(n-1)-connected.
-Corollary $3.$
+Corollary 4.3.
 Let $L$ be a finite flag simplicial complex and let $n\in\mathbb{N}$ with $n\geq 1$.
@@
 Example 4.4.
-Let $L$ be the flag simplicial complex spanned by the graph in Figure 1.
-The right-angled Artin group $A_{L}$ is isomorphic to $F_{2}\times F_{2}\times(\mathbb{Z}^{2}\ast\mathbb{Z})$.
-The group $A_{L}$ is not $(1,2)$-coherent by Theorem 1.5 because $L$ contains a cycle of length 44 as a full subcomplex.
-The group $A_{L}$ is also not $(2,3)$-coherent by Corollary 3.3 because $L$ is 11-connected and not 22-connected.
-On the other hand, the group $A_{L}$ is $(4,\infty)$-coherent by Theorem 1.7 because $L$ is 33-dimensional and hence lies in the class $\mathcal{T}_{4}$.
-We do not know if the group $A_{L}$ is $(3,\infty)$-coherent, although we suspect it to be.
-Note that $L$ does not lie in the class $\mathcal{T}_{3}$, while every full subcomplex of $L$ that is 22-connected is 33-connected.
-$\bullet\bullet\bullet\bullet\bullet\bullet\bullet$
-Figure 1.
-Remark 3.5.
-There is an analogous notion of higher homological coherence [KV23, Fis].
-Let $R$ be a ring and let $n\in\mathbb{N}$.
-A group $G$ is of type $\mathsf{FP}_{n}(R)$ (resp.
-of type $\mathsf{FP}_{\infty}(R)$) if the trivial RGRG-module $R$ admits a projective resolution that is finitely generated in degrees $\leq n$ (resp.
-in all degrees).
-Let $n,m\in\mathbb{N}\cup\{\infty\}$ with $n<m$.
-We say that a group $G$ is homologically $(n,m)$-coherent over $R$ if every subgroup of $G$ that is of type $\mathsf{FP}_{n}(R)$ is of type $\mathsf{FP}_{m}(R)$.
-For $n\geq 2$, homological $(n,m)$-coherence over $\mathbb{Z}$ implies $(n,m)$-coherence because for $n\geq 2$, a group is of type $\mathsf{F}_{n}$ if and only if it is of type $\mathsf{F}_{2}$ and of type $\mathsf{FP}_{n}(\mathbb{Z})$.
-Inheritance results similar to the ones proved above hold also for higher homological coherence.
-Groups of cohomological dimension $n$ over $R$ are homologically $(n,\infty)$-coherent over $R$ [Bro82, Proposition VIII.6.1].
-Over the field $\mathbb{Q}$ (and suitably also over other fields), this can be improved if the top-dimensional $\ell^{2}$-Betti number vanishes [Fis, Proposition 2.3]: If $G$ is a locally indicable group with $\operatorname{cd}_{\mathbb{Q}}(G)=n$ and $b^{(2)}_{n}(G)=0$, then $G$ is homologically $(n-1,\infty)$-coherent over $\mathbb{Q}$.
-Let $L$ be the flag simplicial complex spanned by the graph in Figure 1. The right-angled Artin group $A_{L}$ satisfies $\operatorname{cd}_{\mathbb{Q}}(A_{L})=4$ and by [DL03]
-$b^{(2)}_{4}(A_{L})=\dim_{\mathbb{Q}}H_{3}(L;\mathbb{Q})=0.$
-Hence $A_{L}$ is homologically $(3,\infty)$-coherent over $\mathbb{Q}$ (and, in fact, over every field [FHL24, Proposition 3.15][Fis, Proposition 2.3]).
+Let $L$ be the flag simplicial complex that is the 2-fold suspension of $\Delta^{2}\sqcup\Delta^{0}$.
+Then $L$ is homotopy equivalent to a 2-dimensional sphere.
+The right-angled Artin group $A_{L}$ is isomorphic to $F_{2}\times F_{2}\times(\mathbb{Z}^{3}\ast\mathbb{Z})$.
+The group $A_{L}$ is not $(1,2)$-coherent by Theorem 1.5 because $L$ contains a cycle of length $4$ as a full subcomplex.
+The group $A_{L}$ is also not $(2,3)$-coherent by Corollary 4.3 because $L$ is $1$-connected and not $2$-connected.
+On the other hand, the group $A_{L}$ is $(4,\infty)$-coherent by Theorem 1.7 because $L$ is $4$-dimensional with $H_{4}(L;\mathbb{Q})=0$ and hence lies in the class $\mathcal{T}_{4}$.
+We do not know if the group $A_{L}$ is $(3,\infty)$-coherent.
+Note that $L$ does not lie in the class $\mathcal{T}_{3}$, while every full subcomplex of $L$ that is $2$-connected is $3$-connected.
 Acknowledgements.
-We thank Sam Fisher and Marco Linton for helpful comments. The first author was partially supported by the SFB 1085 Higher Invariants (Universität Regensburg, funded by the DFG).
+We thank Sam Fisher and Marco Linton for very helpful comments. The first author was partially supported by the SFB 1085 Higher Invariants (Universität Regensburg, funded by the DFG).